# Towards Unified Latent Space for 3D Molecular Latent Diffusion Modeling

### v2

@@ 5 Related Works
Molecular Variational Autoencoders. Variational Autoencoders (VAEs) have been widely explored for generative modeling Kingma and Welling (2014); Simonovsky and Komodakis (2018); van den Oord et al. (2017); Razavi et al. (2019), providing a framework to encode data into a structured latent space while enabling both generation and reconstruction Kingma and Welling (2014). In molecular generation, VAEs have been applied to 2D molecular graphs, as seen in JT-VAEJin et al. (2018), MGCVAELee and Min (2022), SSVAEKang and Cho (2019), and CGVAE Liu et al. (2018). However, these methods have several limitations. First, they primarily focus on 2D molecular representations and often rely on non-transformer architectures, limiting their scalability for more complex molecular generation tasks Du et al. (2022). Additionally, VAE-based methods typically assume a simple Gaussian prior, which may fail to accurately model the complex posterior distributions required for effective molecular generation Dai and Wipf (2019); Aneja et al. (2021). With recent advancements in Latent Diffusion Models (LDMs) demonstrating success across multiple domains Vahdat et al. (2021); Zeng et al. (2022), LDMs have been introduced for 3D molecule and protein generation, as seen in GeoLDMXu et al. (2023) and PepGLADKong et al. (2024). These approaches leverage diffusion modeling in latent space, offering improved generative capacity and efficiency.
6 Conclusion and Future Works
In this paper, we propose UAE-3D to compress the multi-modal features of 3D molecules into a unified latent space, and demonstrate the effectiveness of latent diffusion modeling on this space by introducing UDM-3D. By integrating atom types, chemical bonds, and 3D coordinates into a single latent space, our model effectively addresses the inherent challenges of multi-modality and SE(3) equivariance for 3D molecule generation. Extensive experiments on GEOM-Drugs and QM9 confirm leading performances in both de novo and conditional 3D molecule generation, setting new benchmarks for both quality and efficiency.
The success of UAE-3D highlights the benefits of building a unified latent space for molecular design. Moving forward, we will investigate the transferability of the unified latent space to broader molecular design tasks, including structure-based drug design, and extend it to additional modalities, such as proteins, to further assess its effectiveness.
Appendix A Appendix
A.1 Visualization of Representations under Different SE(3) Augmentations
Figure 4 shows the t-SNE visualizations of molecular representations under different SE(3) augmentations. Each dot represents a molecular embedding in the latent space, the colors indicate different molecules, and the color gradient from light to dark corresponds to an increase in the translation distance (left) or the rotation angle (right). Specifically, the left panel (Translation) depicts the embeddings under translation along the same direction, while the right panel (Rotation) illustrates the embeddings under rotation around the same axis.
(a) Translation. Visualization of molecular representations under translation along the same direction. The color gradient indicates increasing translation distance.
(b) Rotation. Visualization of molecular representations under rotation around the same axis with same direction. The color gradient indicates increasing rotation angle.
Figure 4: t-SNE visualizations of molecular representations under different SE(3) augmentations.
A.2 Experimental Details
A.2.1 Evaluation Metrics
@@
Our evaluation framework for de novo 3D molecule generation encompasses complementary metrics to assess molecular validity, diversity, and geometric fidelity. These metrics operate at two levels:
@@
    • Atom Stability: Percentage of atoms with chemically valid valency
    • Validity & Completeness (V&C): Fraction of fully connected, syntactically correct molecules
@@
    • V&U/V&U&N: Rates uniqueness and novelty relative to training data
  • Distribution Alignment: Compares generated/test distributions:
    • FCD: Fréchet distance in ChemNet feature space
    • SNN/Frag/Scaf: Tanimoto similarity for nearest neighbors, BRICS fragments, and Murcko scaffolds
@@
  • Conformer Quality: Evaluates reconstructed 3D coordinates:
    • Atom Stability: Consistency of valency in 3D-derived 2D graphs
@@
    • Bond lengths
    • Bond angles
    • Dihedral angles
A.2.2 Targeted Quantum Properties
For conditional generation, we optimize molecules toward six key electronic properties from quantum chemistry:
@@ A.2.2 Targeted Quantum Properties
The evaluation protocol follows a rigorous split-and-validate strategy Hoogeboom et al. (2022):
@@ A.3 Implement Details
9:     $\mathbf{Z}^{(t-1)}\leftarrow\text{DDPM\_Step}(\mathbf{Z}^{(t)},\tilde{\boldsymbol{\epsilon}},t)$ Ho et al. (2020)
@@
Hyperparameters. We refer insterested readers to our open-source code https://anonymous.4open.science/r/UAE-3D/ for a comprehensive view of our hyperparameters, such as learning rate and hidden dimension.
A.3.1 Algorithm

### v3

@@ Appendix C More Related Works
Molecular Variational Autoencoders. Variational Autoencoders (VAEs) have been widely explored for generative modeling [16, 45, 46, 47], providing a framework to encode data into a structured latent space while enabling both generation and reconstruction [16]. In molecular generation, VAEs have been applied to 2D molecular graphs, as seen in JT-VAE [1], MGCVAE [48], SSVAE [49], and CGVAE [50]. However, these methods have several limitations. First, they primarily focus on 2D molecular representations and often rely on non-transformer architectures, limiting their scalability for more complex molecular generation tasks [51]. Additionally, VAE-based methods typically assume a simple Gaussian prior, which may fail to accurately model the complex posterior distributions required for effective molecular generation [52, 53]. Notably, masked graph modeling [54] also builds molecular a graph auto-encoder, but aiming for representation learning, instead of molecule generation. With recent advancements in Latent Diffusion Models (LDMs) demonstrating success across multiple domains [55, 56], LDMs have been introduced for 3D molecule and protein generation, as seen in GeoLDM[17] and PepGLAD[18]. These approaches leverage diffusion modeling in latent space, offering improved generative capacity and efficiency.
Comparison to Other Unified Generative Molecule Models. Concurrent with our work, ADiT [35] from FAIR also explores latent diffusion for 3D molecules using a variational autoencoder to compress structures into a unified latent space. However, our approach differs in three key ways: (1) ADiT does not achieve near-lossless compression. In contrast, we explicitly evaluate reconstruction quality and show significantly lower errors than ADiT (cf. Table 9), which translates to better 3D generation performance (cf. Table 2). (2) ADiT omits chemical bonds as part of the inputs and reconstruction targets, leading to information loss. Although bond types can be inferred from 3D coordinates, such predictions are not 100% accurate. For example, the QM9 dataset [24] reports 3% inconsistency across inference methods. To address this, we explicitly model bonds using a relational transformer [21] and include bond reconstruction as an objective, offering a more robust and chemically faithful modeling approach. (3) ADiT provides insufficient evaluation for small molecule generation. ADiT reports only two metrics (validity and uniqueness) for the QM9 dataset, while our evaluation includes 15 metrics, covering 2D structure, 3D geometry, stability, validity, uniqueness, and other critical aspects (cf. Table 1). We conjecture that this is because ADiT is more focused on crystal material generation, instead of small molecule generation.
MolCRAFT [41] also aims to unify multiple molecular modalities for generation. However, it processes atom types and coordinates separately—concatenating them while use special design to preserve SE(3) equivariance for coordinates. In contrast, our model encodes atom types, coordinates, and bond types into a single unified latent representation, enabling effortless multi-modal modeling using a uni-modal latent diffusion model – Diffusion Transformer [23]. Moreover, MolCRAFT focuses on the task of structure-based drug design, which is different than our major benchmarks.
Comparison to Other Equivariant 3D Molecule Generation Methods. We adopt SE(3) data augmentations to train a variational autoencoder without explicitly incorporating equivariant architectures, allowing the model to learn geometric symmetries directly from data. This strategy is motivated by AlphaFold3 [57], which demonstrated that SE(3) augmentations are sufficient for general-purpose networks to learn equivariance and invariance. As shown in Table 6, our model achieves near-zero reconstruction errors on 3D molecules subjected to random rotations and translations, showing successful learning of SE(3) equivariance. Compared to models with built-in SE(3) equivariance, such as EDM [19] and JODO [6], our method avoids the architectural complexities of handling multi-modality and enforcing SE(3) equivariant constraints. This design choice allows us to use the highly optimized and parallelizable diffusion model, Diffusion Transformer [23], for latent diffusion modeling. Consequently, we achieve significant efficiency gains, with up to 8.6x faster training and 9.8x faster inference (cf. Figure 2(c;d)).
Appendix D More Experimental Settings
D.1 Experimental Settings for De Novo 3D Molecule Generation
Setup. For the GEOM-Drugs dataset, we use only the ground-state conformer with the lowest energy for each molecule, following the setup in JODO [6]. We adopt the same train/validation/test split ratio of 8:1:1, ensuring a fair comparison. For the QM9 dataset, we use the 100K/10K/10K train/validation/test split. For both datasets, all atoms are shifted to zero center-of-mass before input. All baseline results are either directly borrowed from JODO [6] or reproduced using their released codebase under identical settings.
Hyperparameters. 11 shows the key hyperparameters used for training the UAE-3D and UDM-3D models.
Table 11: Hyperparameters of the UAE-3D and UDM-3D models.
Parameter UAE UDM
epochs 2000 10000
atom loss weight ($\gamma_{\text{atom}}$) 1.0 -
bond loss weight ($\gamma_{\text{bond}}$) 1.0 -
coordinate loss weight ($\gamma_{\text{coordinate}}$) 1.0 -
distance loss weight ($\gamma_{\text{distance}}$) 1.0 -
bonded distance loss weight ($\lambda$) 10.0 -
KLD loss weight 1e-8 -
batch size 512
optimizer AdamW
learning rate 1e-4
weight decay 1e-5
translation augmentation scale 0.1
encoder hidden size 64
encoder #heads 8
encoder #blocks 6
latent dimension 16
decoder hidden size 64
decoder #heads 8
decoder #blocks 4
diffusion hidden size - 512
diffusion #heads - 8
diffusion #layers - 8
diffusion mlp ratio - 4.0
condition drop ($p_{\text{drop}}$) - 0.1
@@ D.1 Experimental Settings for De Novo 3D Molecule Generation
Evaluation Metrics. Our evaluation framework for de novo 3D molecule generation encompasses complementary metrics to assess molecular validity, diversity, and geometric fidelity. These metrics operate at two levels:
@@
    • Atom Stability: Percentage of atoms with chemically valid valency, determined by cross-referencing the bond counts with allowed configurations for each atom type (e.g., carbon with 4 bonds). This is evaluated using RDKit’s atom valency parser. Formal charges are taken into account in the valency specification.
    • Molecule Stability: Percentage of molecules with valid valency for all atoms pass the above valency check.
    • Validity & Completeness (V&C): Fraction of fully connected, syntactically correct (SMILES-parsable) molecules, excluding fragmented or hypervalent structures.
@@
    • Validity & Uniqueness (V&U): Percentage of unique molecules among valid ones, calculated as the ratio of unique SMILES strings to the total number of valid molecules.
    • Validity & Uniqueness & Novelty (V&U&N): Percentage of unique molecules among valid ones that are also novel, calculated as the ratio of unique SMILES strings to the total number of valid molecules, excluding those present in the training set.
  • Distribution Alignment: Convert all the valid molecules into SMILES strings and compare generated/test distributions on the MOSES benchmark [3]:
    • Fréchet ChemNet Distance (FCD): Similarity between generated and reference distributions using activations from the penultimate layer of ChemNet [58], a pretrained molecular property predictor. Lower scores indicate better alignment.
    • Similarity to the nearest neighbor (SNN): Tanimoto similarity between the generated molecules and their nearest neighbors in the test set, calculated using Morgan fingerprints.
    • Fragment Similarity (Frag): Tanimoto similarity between the generated molecules and their nearest neighbors in the test set, calculated using BRICS [59] fragments.
    • Scaffold Similarity (Scaf): Frequencies of Bemis-Murcko scaffolds [60] in the generated molecules compared to the test set.
@@
  • Conformer Quality: The 3D coordinates of generated molecules are converted into bond graphs using a predefined lookup of atomic distance thresholds (per atom pair). Based on these bonds, 2D molecular graphs are reconstructed and then evaluated for atom and molecule stability using the same criteria as in 2D Structural Analysis.
    • Atom Stability: Consistency of valency in 3D-derived 2D graphs. Reconstructs 2D bond orders from 3D atomic distances (via tabulated bond-length thresholds) and checks valency compliance.
    • Molecule Stability: Percentage of molecules with valid valency for all atoms pass the above valency check.
@@
    • Bond lengths: MMD between distributions of bond lengths for eight frequent bond types (e.g., C-C, C-N, C-O). Computed using Gaussian kernel with bandwidth tuned via the median heuristic.
    • Bond angles: MMD between the bond angle distributions of generated vs. reference molecules. Triplets of bonded atoms (e.g., C-C-C) define each angle.
    • Dihedral angles: MMD between distributions of dihedral (torsional) angles over four-atom sequences (e.g., C-C-C-C). Captures conformational realism.
Comparison to Other Evaluation Protocols. Discrepancies in evaluation methodologies can significantly impact reported performance metrics. For instance, ADiT’s validity scores stem from their reliance on pymatgen.core.Molecule for molecular reconstruction which prioritizes topological correctness over chemical stability. While this method validates basic atomic connectivity, it fails to rigorously assess stability and geometric feasibility as evidenced by ADiT’s poor performance on these metrics (Table 2). In contrast, our evaluation employs RDKit’s Chem.Mol for molecule reconstruction, which enforces stricter valency checks and chemical correctness. And our broad evaluation criteria comprehensively measure the validity, diversity, and geometric fidelity of 3D molecule generation.
D.2 Experimental Setting for Conditional 3D Molecule Generation
Setup. We conduct conditional molecule generation on the QM9 dataset following the protocol in EDM [5] and EEGSDE [8]. The properties are evaluated by the QM9 property predictor [24] pre-trained on half of the QM9 dataset. We use the property classifier pre-trained by EDM [5].
Specially, we optimize molecules toward six key electronic properties from quantum chemistry for conditional generation:
@@ D.2 Experimental Setting for Conditional 3D Molecule Generation
The evaluation protocol follows a rigorous split-and-validate strategy [5]:
@@
9:     $\mathbf{Z}^{(t-1)}\leftarrow\text{DDPM\_Step}(\mathbf{Z}^{(t)},\tilde{\boldsymbol{\epsilon}},t)$ [15]
@@
D.3 Pseudo Code